one number is 6 more than another. If the sum of the smaller number and 3 times the larger number is 34, find the two numbers.
Answer by checkley71(8403)

X+6=Y
Y+3(X+6)=34
(X+6)+3(X+6)=34
X+6+3X+18=34
4X=34-6-18
4X=10
X=10/4
X=2.5 ANSWER.
2.5+6=Y
Y=8.5 ANSWER.
PROOF
8.5+3(2.5+6)=34
8.5+3(8.5)=34
8.5+25.5=34
34=34